It took the pilot an hour and a half to make a flight of 240 miles when flying against a headwind. The return trip took an hour and twelve minutes ( the wind had not shifted nor change its speed). What was the speed of the wind?
Answer by lwsshak3(6469)

It took the pilot an hour and a half to make a flight of 240 miles when flying against a headwind. The return trip took an hour and twelve minutes ( the wind had not shifted nor change its speed). What was the speed of the wind?
**
hour and a half=1.5 hr
hour and twelve minutes=1.2 hr
..
let x=wind speed
Flying against the wind:
speed of plane in still air-x=240/1.5
speed of plane in still air=240/1.5+x
..
Flying with the wind:
speed of plane in still air+x=240/1.2
speed of plane in still air=240/1.2-x
..
240/1.5+x=240/1.2-x
2x=240/1.2-240/1.5=.3(240)=72
x=36 mph
ans:
wind speed=36 mph